find the x and y intercepts of the graph of -2x-4y=-20
Answer by drj(1380)

Find the x and y intercepts of the graph of -2x-4y=-20

Step 1. The intercepts occur when the one of the coordinates is zero. The y-intercept is when x=0 or at point (0,b) where b is the y-intercept. The x-intercept is when y=0 or at point (a,0) where a is the x-intercept.

Step 2. So let x=0, then -2*0-4y-20 or 4y=20 or y=5 after dividing by 4 to both sides. Therefore, the y-intercept is 5.

Step 3. So let y=0, then -2x-4*0-20 or 2x=20 or x=10 after dividing by 10 to both sides. Therefore, the x-intercept is 10.

Step 4. ANSWER: The x-intercept is 10 and the y-intercept is 5.

I hope the above steps and explanation were helpful.

For Step-By-Step videos on Introduction to Algebra, please visit http://www.FreedomUniversity.TV/courses/IntroAlgebra and for Trigonometry please visit http://www.FreedomUniversity.TV/courses/Trigonometry.

Also, good luck in your studies and contact me at john@e-liteworks.com for your future math needs.

Respectfully,
Dr J

http://www.FreedomUniversity.TV